### v2

# P-ICL: Point In-Context Learning for Named Entity Recognition with Large Language Models

Abstract
In recent years, the rise of large language models (LLMs) has made it possible to directly achieve named entity recognition (NER) without any demonstration samples or only using a few samples through in-context learning (ICL). However, standard ICL only helps LLMs understand task instructions, format and input-label mapping, but neglects the particularity of the NER task itself. In this paper, we propose a new prompting framework P-ICL to better achieve NER with LLMs, in which some point entities are leveraged as the auxiliary information to recognize each entity type. With such significant information, the LLM can achieve entity classification more precisely. To obtain optimal point entities for prompting LLMs, we also proposed a point entity selection method based on K-Means clustering. Our extensive experiments on some representative NER benchmarks verify the effectiveness of our proposed strategies in P-ICL and point entity selection.111Our code is available at https://github.com/jiangguochaoGG/P-ICL.
P-ICL: Point In-Context Learning for Named Entity Recognition
with Large Language Models
Guochao Jiang†, Zepeng Ding†, Yuchen Shi†, Deqing Yang †School of Data Science, Fudan University, Shanghai, China ‡Shanghai Key Laboratory of Data Science, Shanghai, China †{gcjiang22, zpding22, ycshi21}@m.fudan.edu.cn ‡yangdeqing@fudan.edu.cn
1 Introduction
Named Entity Recognition (NER) Li et al. (2022) is a critical task of information extraction and plays a vital role in various downstream applications, including knowledge graph Xu et al. (2017) construction, improving information retrieval systems Banerjee et al. (2019) and question-answering systems Aliod et al. (2006). The primary goal of NER is to precisely identify and classify the spans of entity mentions within a sentence into their corresponding entity types.
Large Language Models (LLMs), such as ChatGPT222https://openai.com/blog/chatgpt and GPT-4 OpenAI (2023), have introduced new research directions in various natural language processing tasks. This shift is largely due to the in-context learning (ICL) abilities of LLMs, allowing them to perform tasks with only a few examples. With the rapid progress of LLMs, exploring their direct use or applying few-shot learning techniques to NER tasks is highly valuable. This strategy, which utilizes LLMs without extensive fine-tuning, offers a potential way to significantly lower the training costs associated with NER tasks.
Existing research has often neglected the unique aspects of using LLMs for NER through ICL, especially when compared to their usage in other natural language processing (NLP) tasks. Previous study Min et al. (2022) has shown that improvements in LLMs’ performance through ICL mainly come from input-label mapping and format following. Some recent works Chen et al. (2023); Lee et al. (2022); Wang et al. (2023b, c) have explored the use of ICL for NER, but they ignored the role of entity examples explored in this paper. In fact, for NER, the information of entity type, along with entity classification, are the key factors to achieve overall task success. Unfortunately, such information has not yet been emphasized in previous ICL-based approaches.
To tackle this issue, we propose a new framework for NER in this paper, named P-ICL (Point In-Context Learning). Our framework equips the LLM with some representative entity instances for each entity type in the prompt, which provides significant information on entity type and entity classification. Compared with the demonstration examples in standard ICL’s prompt, the representative entity instances added in P-ICL’s prompt can be regarded as the data points in entity space, which are denoted as Point entities in this paper. As shown in Figure 1, providing these point entities for each entity type in the prompt greatly helps the LLM accurately understand the semantic distinctions between different entity types and their related domains. This enhancement is crucial for improving the NER performance of LLMs. Furthermore, designing effective point entity selection methods to select more representative point entities can help LLMs obtain entity type information and entity classification more effectively.
Our main contributions in this paper are summarized as follows.
1.
We propose a novel framework P-ICL for LLMs to achieve NER. Our framework leverages point entities to provide the LLM with the significant information of entity type and entity classification, that is overlooked by standard ICL.
2.
The point entity selection method we proposed achieves better performance in the experiments compared with random selection, illustrating the significance of selecting more representative point entities for enhancing P-ICL’s capability of NER.
3.
Our results of using three LLMs upon six representative NER benchmarks demonstrate the advantages of our proposed P-ICL over the standard ICL method.
2 Related Work
NER is an important task in the field of information extraction, which focuses on identifying named entities of interest from sentence-level or document-level text Sang and Meulder (2003). The existing NER solutions can be divided into three main categories: sequence labeling, span classification and generative methods. Using the sequence labeling approaches Ratinov and Roth (2009); Straková et al. (2019); Dai et al. (2020) split the text into tokens and the model assigns a label like BIO tagging to each token. Methods that use span classification Wang et al. (2020); Yu et al. (2020) take the text into account in the form of the basic elements of a span, and then for the span decide whether it is an entity. Generative methods, on the other hand, rely on the overall framework of generative modeling to output the results of the final NER in the form of text generation Cui et al. (2021); Yan et al. (2021).
In recent years, LLMs have shown promising results on many NLP tasks including NER. Wei et al. (2023b) explored the direct use of ChatGPT for zero-shot named entity recognition, relation extraction and event extraction, and proposed a framework named ChatIE to interact with LLMs to accomplish the task of information extraction. Similarly, Ji (2023) explored the performance of the open-source model Vicuna Zheng et al. (2023) on NER in the zero-shot and few-shot scenarios. Hu et al. (2023) explored the NER capabilities of ChatGPT within a specific clinical domain, revealing the possibilities of applying generalized language models in specialized domains. Xie et al. (2023a) presented a systematic empirical investigation of zero-shot NER with LLM and adapted prevalent reasoning methods to NER. Xie et al. (2023b) explored the possibility of boosting zero-shot NER with LLMs via self-improving. Jiang et al. (2024b) introduced a generative NER framework named ToNER with entity type matching, which shows that entity mention is conducive to further improving NER performance. GoLLIE Sainz et al. (2023) employed an input-output format that includes schema definition, input text, and output annotations to perform IE tasks, but it ignored the role of point entity in this paper and the selection method for data examples.
In-context learning (ICL), as an emergent capability of LLMs, allows LLMs to quickly have the ability to adapt to downstream tasks given a few task examples in the prompt. The work that has been done focuses on the working mechanism of ICL and the factors that influence it Min et al. (2022); Lyu et al. (2023); Wei et al. (2023a). Min et al. (2022) found that the label space of the examples, the distribution of the input text and the overall form of the sequence were factors that influenced ICL performance. Wang et al. (2023a) delves into the mechanism of label words in ICL from the information flow perspective and introduces a re-weighting method to enhance the ICL performance of LLM based on this finding.
3 Preliminaries
NER aims to extract entities from the given text and assign the correct entity type to the entities. Formally, we denote the input token sequence as . With LLM , we denote the output token sequence as under the instruction . Instruction usually includes the entity types that the task focuses on such as PER, LOC, ORG and MISC in the CoNLL2003 dataset Sang and Meulder (2003).
LLMs can learn from demonstrations to improve the downstream performance with ICL ability. For NER, represents the input text, represents the extraction result, and represents the number of demonstrations. Given some demonstrations , LLMs can output results that are more consistent with the task instructions and these examples. Therefore, in the case of ICL, the extraction result of model for text is
To facilitate the analysis of the generated results, we stipulate that must be a parsable JSON result, where the key represents the entity type and the value is the entity list of the entity type corresponding to the key.
4 Methodology
In this section, we initially discuss the standard ICL and point entity. Subsequently, we introduce our P-ICL and the method for selecting point entities. Figure 2 illustrates the overview of our P-ICL pipeline, including the point entity selection process.
4.1 Standard ICL
In the previous discussion, we introduced a JSON schema aimed at facilitating the generation of structured and easily parsable outputs by LLM. Nonetheless, it is crucial to emphasize that without employing prompt engineering techniques, such as ICL, the outputs from LLM may not consistently adhere to the specified requirements.
ICL has been shown through empirical research to help LLMs learn the necessary input-output mapping relevant to specific downstream tasks. This strategy plays a key role in minimizing instances where LLM outputs fail to meet task requirements. In this study, which concentrates on the NER task, we employ the CoNLL2003 dataset as a case in point. We detail and apply particular prompts to carry out zero-shot and few-shot experiments under the ICL framework.
Zero-shot:
Please list all named entities of the following entity types in the input sentence:
- PER
- ORG
- LOC
- MISC
You should output your results in the format {"type": [entity]} as a json.
Input:
Few-shot with ICL:
Please list all named entities of the following entity types in the input sentence:
- PER
- ORG
- LOC
- MISC
Here are some examples:
Input:
Output:
Input:
Output:
You should output your results in the format {"type": [entity]} as a json.
Input:
4.2 Point Entity
In the approach described, the standard ICL method requires a certain number of examples to understand the details of downstream tasks. However, it does not fully capture the unique features of the NER task. As we know, two main factors affect NER performance: contextual text and entity mention. Typically, the contextual text provides extensive semantic information about a sentence, while the entity mention clearly identifies the potential entity. For LLMs, pre-training on a wide corpus gives the model a comprehensive understanding of natural language. As a result, many high-quality LLMs have strong text comprehension abilities. This is why LLMs perform well in various NLP tasks, even without specific training or in a few-shot ICL scenario.
Entity mention, referring to how natural language identifies actual named entities, should be a key focus of NER for LLMs aiming for advanced text understanding. From a human intuition standpoint, entity mentions are usually quite specific. A major challenge in NER involves effectively defining the scope of each entity type in the downstream tasks. For LLMs, the strategies to address this challenge generally fall into two categories: providing detailed and clear definitions or presenting examples of entities. Embedding a specific number of entity instances for each entity type in the prompt can help LLMs capability discriminate the boundaries of relevant entity types, thereby boosting LLMs’ performance on NER. Since this method only requires the presentations of entity mentions, which aligns well with the concept of a data point, we denote it as Point entity.
4.3 P-ICL: Point ICL
The concept of a point entity offers a novel approach to prompting LLM for NER task. By specifying the entities of interest through sufficient point entities, we can effectively communicate the types of entities to the LLM. This method, when compared to the standard ICL technique, allows for the inclusion of more training data for the LLM. Unlike standard ICL, which necessitates the full context, the point entity approach only requires the entity mention itself.
The ICL-like method that provides data information to LLM based on point entities is called P-ICL. Formally, assume that the entity type label set of the NER task is , and the number of entity types considered is . For the entity type , there are point entities, respectively . In this paper, for the NER task, taking the CoNLL2003 dataset as an example, we consider the following prompts to conduct experiments under P-ICL:
P-ICL:
Please list all named entities of the following entity types in the input sentence:
- PER: e.g. , ,,
- ORG: e.g. , ,,
- LOC: e.g. , ,,
- MISC: e.g. , ,,
Here are some examples:
Input:
Output:
Input:
Output:
You should output your results in the format {"type": [entity]} as a json.
Input:
4.4 Point Entity Selection
Building on the previous discussion, it’s clear that the role of point entities is to provide LLMs with insights about different types of entities through mentions. Therefore, it makes sense to carefully select the right point entity. When point entities effectively represent the distribution of the entire dataset, LLMs can use their advanced natural language understanding to gain more information from inputs that are more representative, compared to when point entities are chosen randomly. This highlights the importance of careful selection of point entities to improve the performance of LLMs in processing and interpreting data, thus maximizing the information obtained from the inputs.
Recent advancements in Pre-trained Language Models (PLMs) like BERT Devlin et al. (2019) have provided many solutions for NLP tasks through representation learning. This approach mainly uses PLMs to extract relevant features from text for use in various tasks. In a given dataset, all the data in the training set can be used to help select representative entities. Since each entity can have a corresponding representation, choosing some representative entities aligns with unsupervised learning methods. In this scenario, the use of clustering algorithms Ezugwu et al. (2022) proves to be an extremely effective technique.
By using PLMs alongside common clustering algorithms, it’s possible to encode training entities related to a specific entity type to obtain their representations. Then, by clustering these representations, one can pinpoint central entities. The selection of a central entity, based on its position in the representation space, naturally reflects its representativeness. This method highlights the benefit of combining PLMs with clustering algorithms for a refined identification and selection of central entities. This enhances the model’s capacity to generalize from representative examples in unsupervised settings.
Formally, let be the entity set for entity type . For each entity of entity type , we can use the following method to obtain its corresponding representation:
where is the representation dimension, which generally depends on the encoder used. Let be the corresponding entity representation set. Consistent with the above discussion, for entity type , we use the following form to obtain entity centers with classic K-Means clustering algorithm:
Then we need to convert the corresponding representation in the representation space into a string entity that can be used in prompt. For each entity center , we obtain the corresponding point entity by maximizing the similarity between the representation of the real entity and its entity center:
where the function measures the degree of similarity between two given elements, and is generally chosen as cosine similarity in most NLP tasks.
5 Experiments
In this section, we will evaluate the results of our proposed P-ICL method on some NER datasets, using different closed-source and open-source models. 333Implementation details are in Appendix A
5.1 Datasets
Dataset | Training | Training | Test | Test |
Name | Sen. # | Ent. # | Sen. # | Ent. # |
CoNLL2003 | 14,041 | 8,082 | 3,453 | 2,637 |
WNUT2017 | 3,394 | 703 | 1,287 | 344 |
ACE2004 | 6,202 | 10,090 | 812 | 1,692 |
ACE2005 | 7,299 | 9,873 | 1,060 | 1,573 |
JNLPBA | 18,545 | 18,899 | 3,856 | 4,344 |
BC5CDR | 5,228 | 2,719 | 5,865 | 2,713 |
We conducted our experiments on the following NER benchmarks. These datasets’ statistics are listed in Table 1.
CoNLL2003 CoNLL2003 Sang and Meulder (2003) is a collection of news wire articles from the Reuters Corpus, which contains 4 entity types including LOC, ORG, PER and MISC.
WNUT2017 WNUT2017 Derczynski et al. (2017) focuses on unusual and previously-unseen entities in the context of emerging discussions, which contains 6 entity types including person, location, corporation, product, creative-work and group.
ACE2004 and ACE2005 ACE2004 Doddington et al. (2004) and ACE2005 Walker et al. (2006) are two nested named entity recognition datasets, which contain 7 entity types including person, location, organization, geographical social political, weapon, facility and vehicle. We followed the same dataset split setup as the previous work Katiyar and Cardie (2018); Lin et al. (2019).
JNLPBA JNLPBA Collier and Kim (2004) is a biomedical NER dataset, which contains 5 entity types including DNA, protein, cell_type, cell_line and RNA.
BC5CDR BC5CDR Li et al. (2016) is a large annotated text corpus of human annotations of all chemicals, diseases and their interactions, which contains 2 entity types including Chemical and Disease.
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | JNLPBA | BC5CDR | Avg. | |
Vanilla | 52.47 | 11.96 | 17.06 | 14.35 | 34.46 | 72.18 | 33.75 | |
Standard ICL | 3-shot | 65.15 | 13.49 | 32.27 | 32.85 | 45.06 | 71.54 | 43.39 |
5-shot | 67.24 | 14.21 | 36.70 | 37.10 | 46.77 | 72.07 | 45.68 | |
10-shot | 71.01 | 14.68 | 43.45 | 43.50 | 49.25 | 72.24 | 49.02 | |
20-shot | 73.57 | 16.68 | 46.69 | 46.41 | 50.25 | 72.35 | 50.99 | |
P-ICL | 5+3-shot | 70.96 | 16.61 | 37.55 | 39.12 | 50.54 | 72.40 | 47.83 |
10+3-shot | 71.26 | 16.93 | 40.35 | 37.91 | 50.15 | 72.72 | 48.22 | |
5+5-shot | 72.80 | 16.94 | 41.58 | 42.16 | 51.31 | 73.12 | 49.65 | |
10+5-shot | 72.80 | 18.77 | 42.81 | 41.68 | 51.46 | 73.55 | 50.18 | |
5+10-shot | 74.17 | 18.99 | 47.03 | 45.87 | 51.73 | 73.67 | 51.91 | |
10+10-shot | 74.63 | 19.59 | 47.86 | 47.67 | 52.45 | 73.73 | 52.66 | |
5+20-shot | 76.18 | 20.74 | 50.43 | 49.28 | 53.14 | 74.10 | 53.98 | |
10+20-shot | 76.07 | 21.06 | 50.50 | 50.33 | 53.61 | 74.62 | 54.37 | |
P-ICL | 5+3-shot | 75.48 | 22.10 | 43.73 | 42.56 | 53.61 | 77.40 | 52.48 |
10+3-shot | 76.43 | 21.28 | 44.91 | 42.56 | 53.92 | 76.52 | 52.60 | |
5+5-shot | 76.36 | 22.02 | 46.47 | 46.24 | 55.07 | 78.33 | 54.08 | |
10+5-shot | 76.70 | 23.19 | 47.70 | 46.44 | 54.96 | 77.35 | 54.39 | |
5+10-shot | 78.95 | 21.51 | 52.11 | 52.39 | 56.70 | 78.46 | 56.69 | |
10+10-shot | 79.32 | 22.63 | 52.32 | 51.90 | 56.47 | 77.82 | 56.74 | |
5+20-shot | 80.67 | 24.54 | 54.45 | 54.13 | 56.91 | 78.49 | 58.20 | |
10+20-shot | 80.69 | 25.00 | 55.37 | 53.61 | 56.86 | 77.88 | 58.24 |
5.2 Overall Performance
We experimented on the six datasets with a variety of backbone LLMs, including LLaMA-3-70B, GPT-3.5 Turbo, and Mixtral 8x7B. The results for LLaMA-3-70B are documented in Table 2, while the outcomes for the other two models are detailed in Appendix B. For each model, we employed different approaches, ranging from a zero-shot version without demonstrations, referred to as Vanilla, to few-shot versions, labeled ICL, which incorporate a range of example samples. For the P-ICL method proposed in this paper, we tested various selection techniques, including random selection and K-Means clustering-based selection, designated as P-ICL and P-ICL, respectively. For the ICL configurations, we evaluated the model’s performance with different numbers of samples , specifically . Moreover, in the P-ICL models, we experimented with the number of clusters per type, exploring . For instance, in the 5+3-shot experiment for P-ICL, this denotes the selection of five entity examples per type and three NER samples for few-shot demonstrations. Based on the experimental results, we have the following observations:
1.
In the case of Standard ICL, we observed that the model’s results improve with an increasing number of demonstrations until a certain threshold is reached. For example, there is no significant performance enhancement when scaling from 10-shot to 20-shot.
2.
When comparing Standard ICL with P-ICL at equal numbers of demonstrations, the performance of P-ICL is consistently superior. This finding underscores the significance of introducing point entities for the NER task.
3.
Comparing P-ICL with P-ICL under the same setting, the latter outperforms the former. This result validates the effectiveness of our proposed point entity selection method based on K-Means clustering in enhancing the performance of P-ICL.
Furthermore, a comparison of experimental outcomes with and reveals minimal differences, hence, in the following section, we delve further into the analysis of the impact of point entity quantity.
5.3 Effect of Point Entity Supervision
ICL shot | CoNLL2003 | WNUT2017 |
3 | 71.21 | 16.91 |
5 | 72.95 | 17.54 |
10 | 74.02 | 19.33 |
20 | 76.03 | 21.57 |
The source of point entities influences whether the model receives additional supervision beyond ICL. To investigate the impact of point entity supervision, we examine the performance of LLaMA-3-70B under point entity restriction on the CoNLL2003 and WNUT2017 datasets. Specifically, point entities can only originate from ICL example labels, as shown in Table 3. Compared to standard ICL results, the restricted P-ICL results consistently outperform standard ICL performance, demonstrating the advantage of P-ICL even without additional supervision. Additionally, when compared to P-ICL with the same number of ICL shots, the restricted performance shows minimal difference. This further indicates that LLMs can efficiently learn task information from point entities, which will be further discussed in Section 5.4.
5.4 Effect of Point Entity Number
To delve deeper into how the quantity of point entities influences the performance of P-ICL, we conducted experiments on the CoNLL2003 dataset, with the results of LLaMA-3-70B displayed in Figure 3. For a fixed sample size , we varied the number of point entities selected by BERT-large from each category, , ranging from 1 to 10. The findings indicate that: (1) P-ICL outperforms standard ICL when comparing results under an equal number of ICL demonstrations. (2) With the same number of ICL demonstrations, P-ICL performance initially improves as the number of point entities increases. However, when the number of each type of point entity is high, the model’s performance plateaus. This suggests that LLMs can efficiently learn task-relevant information from a given set of point entities, and a smaller number of point entities is sufficient.
5.5 Effect of Point Entity Selection
BERT-large | RoBERTa-large | BGE-large | mxbai-embed-large | |
# Param. | 336M | 355M | 335M | 335M |
5+3-shot | 75.48 | 75.27 | 76.86 | 75.55 |
10+3-shot | 76.43 | 76.67 | 77.42 | 76.07 |
5+5-shot | 76.36 | 77.21 | 78.21 | 77.46 |
10+5-shot | 76.70 | 77.80 | 78.86 | 78.28 |
5+10-shot | 78.95 | 78.66 | 79.98 | 79.30 |
10+10-shot | 79.32 | 79.29 | 80.20 | 79.65 |
5+20-shot | 80.67 | 81.01 | 81.34 | 80.82 |
10+20-shot | 80.69 | 80.44 | 81.29 | 81.02 |
Avg. | 78.06 | 78.29 | 79.27 | 78.52 |
To investigate the influence of point entity selection on performance outcomes, this section presents two sets of experiments. Firstly, we explore the impact of text encoders on results by utilizing different textual encoders for entity representation. Secondly, we examine the effects of varying the number of selected point entities for each entity center.
To examine the effects of different encoders, we selected four representative models: two classic BERT-like models (BERT-large Devlin et al. (2019) and RoBERTa-large Liu et al. (2019)), and two recent retrieval embedding models (BGE-large Xiao et al. (2023) and mxbai-embed-large444https://www.mixedbread.ai/blog/mxbai-embed-large-v1). All of these models have a similar number of parameters. Employing these encoders for entity representation and performing point entity selection based on the methodology described in Section 4.4, we conducted experiments on the CoNLL2003 dataset. The results of LLaMA-3-70B, as depicted in Table 4, indicate that: (1) Utilizing BGE-large and mxbai-embed-large as encoder consistently outperforms BERT-like models, demonstrating that retrieval embedding models can yield more effective point entity selection results. (2) The outcomes for both BERT-like models are comparable, as are the results for the two retrieval embedding models. Overall, BGE-large exhibits superior performance to mxbai-embed-large, hence we have selected BGE-large as the encoder for our final model.
P-ICLτ | P-ICL + Extend | |
1 | 78.95 | 78.95 |
2 | 79.56 | 79.40 |
3 | 79.84 | 79.69 |
4 | 79.57 | 79.85 |
5 | 79.58 | 78.99 |
Avg. | 79.50 | 79.38 |
In Equation 4, we select point entities by maximizing the similarity between the real entity’s representation and the entity center. This approach is predicated on the notion that each entity center requires only one most representative entity. To validate this concept, we experimented with selecting a varying number of point entities for the same entity center, referred to as . When , a single entity center corresponds to multiple point entities. Additionally, to eliminate the influence of the number of point entities in comparative experiments, we expanded the number of point entities by randomly selecting entities post-selection based on Equation 4.
The experimental outcomes of LLaMA-3-70B for a 5+10-shot setting on the CoNLL2003 dataset are presented in Table 5, where P-ICLτ denotes the model selecting most similar entities as point entities for each entity center, and the corresponding P-ICL + Extend represents the model extending P-ICL with additional entities. The results reveal that: (1) Comparing P-ICLτ for and , selecting a larger does not lead to much better outcomes and may even result in a decline, suggesting that each entity center needs only the most representative entity, and selecting more similar entities could lead to redundancy due to excessive similarity. (2) There is no significant difference between the results of P-ICLτ and P-ICL + Extend, implying that compared to point entities obtained through K-Means clustering, selecting more entities for each entity center or employing random extension does not yield better results.
6 Conclusion
In this paper, we introduce P-ICL, a framework that employs point entities to improve LLM in performing NER through ICL. By incorporating point entities, we equip LLMs with specific entity type information and entity classification, addressing the limitations of the standard ICL approach for NER. Our findings, based on six widely used NER datasets and experiments conducted with three renowned LLMs, confirm the efficacy of P-ICL.
Ethics Statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Human annotations are only used in methodological research at the beginning of the work, to assist in analyzing the feasibility of the proposed solution. Annotators consented to the use of data for research purposes. We ensure that the privacy of all annotators is protected throughout the annotation process, and all of them are adequately paid according to local standards. Human annotations are not applied during the evaluation of our method.
Risks
In this paper, all datasets are obtained from official sources. The datasets adopted have been anonymized and do not contain offensive information. However, we cannot guarantee that the datasets do not contain socially harmful or toxic language.
Limitations
We focus solely on evaluating the P-ICL framework’s performance with fixed point entities. For LLMs, having fixed point entities may not always be ideal for varying inputs. This limitation highlights the potential for future work on dynamically choosing point entities depending on the input sentence.
Appendix A Implementation Details
We report the entity-level micro F1 scores in the following result tables and figures to compare the NER performance. To more comprehensively compare the P-ICL performance of existing LLMs, we selected three representative models at this stage, including GPT-3.5 Turbo, LLaMA-3-70B AI@Meta (2024) and Mixtral 8x7B Jiang et al. (2024a). The demonstrations used in ICL are selected randomly in the training set. For standard ICL and P-ICL experiments, due to a certain degree of randomness, the corresponding experimental results are the average of the F1 scores of 5 experiments. For experiments on open-source models, we use the vLLM Kwon et al. (2023) framework for inference. In order to maintain the stability and consistency of model generation, we use the greedy generation to control the outputs of LLMs. We conducted our experiments on eight NVIDIA A800 80GB PCIe GPUs.
Appendix B Overall Performance
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | JNLPBA | BC5CDR | Avg. | |
Vanilla | 62.49 | 8.70 | 17.99 | 14.35 | 38.83 | 68.30 | 35.11 | |
Standard ICL | 3-shot | 64.95 | 10.55 | 20.30 | 19.84 | 42.88 | 67.47 | 37.67 |
5-shot | 66.47 | 10.92 | 22.48 | 22.32 | 43.53 | 68.17 | 38.98 | |
10-shot | 67.65 | 11.51 | 23.94 | 25.42 | 46.09 | 68.67 | 40.55 | |
20-shot | 66.35 | 10.39 | 23.90 | 27.05 | 45.05 | 69.12 | 40.31 | |
P-ICL | 5+3-shot | 66.80 | 16.61 | 24.31 | 24.57 | 45.55 | 67.10 | 40.82 |
10+3-shot | 65.70 | 14.93 | 23.59 | 24.19 | 45.38 | 66.41 | 40.03 | |
5+5-shot | 67.56 | 16.95 | 25.49 | 27.01 | 46.83 | 67.64 | 41.91 | |
10+5-shot | 66.58 | 16.10 | 25.85 | 26.51 | 46.05 | 67.40 | 41.42 | |
5+10-shot | 67.66 | 16.83 | 28.18 | 28.95 | 48.00 | 69.10 | 43.12 | |
10+10-shot | 68.27 | 18.51 | 26.64 | 30.02 | 48.01 | 68.33 | 43.30 | |
5+20-shot | 69.31 | 19.36 | 29.10 | 32.05 | 48.99 | 69.60 | 44.73 | |
10+20-shot | 69.58 | 18.40 | 28.31 | 31.99 | 49.16 | 69.28 | 44.45 | |
P-ICL | 5+3-shot | 68.87 | 19.10 | 26.34 | 25.57 | 48.17 | 69.73 | 42.96 |
10+3-shot | 69.47 | 17.99 | 26.88 | 25.36 | 47.85 | 68.34 | 42.65 | |
5+5-shot | 70.33 | 16.70 | 27.23 | 29.12 | 48.39 | 70.42 | 43.70 | |
10+5-shot | 70.83 | 19.41 | 26.74 | 28.09 | 48.65 | 69.18 | 43.82 | |
5+10-shot | 71.04 | 18.18 | 29.66 | 31.03 | 50.22 | 71.27 | 45.23 | |
10+10-shot | 71.12 | 20.98 | 29.11 | 30.45 | 51.04 | 70.36 | 45.51 | |
5+20-shot | 72.58 | 21.09 | 31.17 | 32.69 | 51.06 | 71.26 | 46.64 | |
10+20-shot | 71.80 | 20.23 | 30.53 | 32.19 | 50.86 | 71.10 | 46.12 |
Method | CoNLL2003 | WNUT2017 | ACE2004 | ACE2005 | JNLPBA | BC5CDR | Avg. | |
Vanilla | 37.38 | 6.31 | 19.01 | 17.26 | 21.85 | 54.63 | 26.07 | |
Standard ICL | 3-shot | 45.93 | 8.76 | 28.22 | 25.42 | 26.95 | 57.69 | 32.16 |
5-shot | 48.10 | 10.11 | 30.54 | 27.90 | 28.16 | 57.68 | 33.75 | |
10-shot | 51.00 | 12.56 | 32.42 | 30.94 | 29.38 | 57.83 | 35.69 | |
20-shot | 51.12 | 11.37 | 32.93 | 29.78 | 32.08 | 56.38 | 35.61 | |
P-ICL | 5+3-shot | 46.37 | 10.62 | 31.09 | 28.79 | 25.89 | 59.11 | 33.65 |
10+3-shot | 45.85 | 11.00 | 30.74 | 28.91 | 25.46 | 56.91 | 33.15 | |
5+5-shot | 48.87 | 11.67 | 31.13 | 31.33 | 28.15 | 58.84 | 35.00 | |
10+5-shot | 48.53 | 12.64 | 33.26 | 31.06 | 27.88 | 58.30 | 35.28 | |
5+10-shot | 53.09 | 13.49 | 35.92 | 34.57 | 29.86 | 59.45 | 37.73 | |
10+10-shot | 53.07 | 13.09 | 35.37 | 35.04 | 31.33 | 57.96 | 37.64 | |
5+20-shot | 56.50 | 14.85 | 37.02 | 36.69 | 32.08 | 58.28 | 39.24 | |
10+20-shot | 56.67 | 14.26 | 36.01 | 35.22 | 30.44 | 57.99 | 38.43 | |
P-ICL | 5+3-shot | 50.59 | 12.54 | 32.18 | 31.27 | 27.87 | 62.14 | 36.10 |
10+3-shot | 48.59 | 13.00 | 33.67 | 30.69 | 29.38 | 58.17 | 35.58 | |
5+5-shot | 51.81 | 14.55 | 36.08 | 34.03 | 29.21 | 61.36 | 37.84 | |
10+5-shot | 50.02 | 13.75 | 37.12 | 33.59 | 31.89 | 59.23 | 37.60 | |
5+10-shot | 56.18 | 15.36 | 36.77 | 36.78 | 32.40 | 61.10 | 39.77 | |
10+10-shot | 54.42 | 15.06 | 37.82 | 36.14 | 32.99 | 59.79 | 39.37 | |
5+20-shot | 59.12 | 16.92 | 39.38 | 38.60 | 35.05 | 60.41 | 41.58 | |
10+20-shot | 58.70 | 16.19 | 41.18 | 38.34 | 32.26 | 59.18 | 40.98 |